Could someone please help me to solve the following system of equations and find the ordered pair?
x+3y=6 (1)
x=7-3y (2)

Answer by ewatrrr(24785)

 
Hi
x = 7-3y
x+3y=6
(7-3y) + 3y = 6 |substituting for x
7 ≠ 6 There is NO solution for this system of equations